Seafoam:

Which of the following could be used to calculate the area of the sector in the circle shown above? π(5in)30 over 360 π(5in)230 over 360 π(30in)25 over 360 π(30in)5 over 360

Seafoam:

1 attachment
Seafoam:

@snowflake0531

snowflake0531:

If I did this I would do 30/360 = x/area and the area is pi r^2, which is 25pi 30/360 = x/25pi and then cross multiply

snowflake0531:

So I got 30/360 because this is looking at degrees so ---? 30/360 But then we also have x/area, because that's the fraction of it we want So x/25pi = 30/360 Cross multiply you would get 360x = 25pi (30) So x equals...?

AZ:

Cross multiplying makes it seem more complicated Just use the formula 

AZ:

\(\text{area of sector }= \dfrac{\theta}{360} \times \pi r^2\) plug in your angle in place of \(\theta\) and replace r with your radius ta-da

Seafoam:

So it has to be one of the last two options?

Seafoam:

Oh wait, now I've just confused myself. Is it switched around

AZ:

what is the radius? what is the angle?

Seafoam:

Oh no I think in the picture you sent the pi r squared was on the right and mine is on the left whoops

Seafoam:

Oh the second one

Seafoam:

Again?

AZ:

it doesn't matter which order, you're multiplying it all it just matters that you make sure you're doing pi * angle/ 360 * radius ^2

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Seafoam Oh the second one \(\color{#0cbb34}{\text{End of Quote}}\) Correct

Seafoam:

yayy

Seafoam:

Thank you!

AZ:

You're welcome!